# Music-Aligned Holistic 3D Dance Generation via Hierarchical Motion Modeling

### v1

@@ Abstract
Well-coordinated, music-aligned holistic dance enhances emotional expressiveness and audience engagement. However, generating such dances remains challenging due to the scarcity of holistic 3D dance datasets, the difficulty of achieving cross-modal alignment between music and dance, and the complexity of modeling interdependent motion across the body, hands, and face. To address these challenges, we introduce SoulDance, a high-precision music-dance paired dataset captured via professional motion capture systems, featuring meticulously annotated holistic dance movements. Building on this dataset, we propose SoulNet, a framework designed to generate music-aligned, kinematically coordinated holistic dance sequences. SoulNet consists of three principal components: (1) Hierarchical Residual Vector Quantization, which models complex, fine-grained motion dependencies across the body, hands, and face; (2) Music-Aligned Generative Model, which composes these hierarchical motion units into expressive and coordinated holistic dance; (3) Music-Motion Retrieval Module, a pre-trained cross-modal model that functions as a music-dance alignment prior, ensuring temporal synchronization and semantic coherence between generated dance and input music throughout the generation process. Extensive experiments demonstrate that SoulNet significantly surpasses existing approaches in generating high-quality, music-coordinated, and well-aligned holistic 3D dance sequences. Additional resources are available at: https://xjli360.github.io/SoulDance/.
@@ 1 Introduction
However, despite recent progress in this field, generating well-coordinated holistic dance sequences that aligned with music remains a challenging task due to three critical limitations. The first constraint refers to dataset inadequacies in which existing music-dance datasets either lack crucial components of dance performance or suffer from quality issues. While datasets like AIST++[31] provide body-only movements with musical correspondence, and FineDance[32] and Choreomaster [7] incorporate fine-grained finger movements, they omit facial expressions—a vital element for conveying emotion in dance. Furthermore, most datasets rely on pose estimation from videos, introducing inherent inaccuracies that compromise data quality. The second limitation concerns modeling constraints where current approaches fail to generate coordinated dance movements due to insufficient capacity for capturing complex hierarchical interdependencies between body, hands and face. The third limitation involves inadequate cross-modal alignment between music and dance, preventing coherent choreographic sequences that consistently match input music.
To overcome these challenges, we present a comprehensive solution encompassing both dataset and methodological innovations. We introduce SoulDance, a high-quality dataset comprising 12.5 hours of paired music and dance data captured using professional motion capture systems. Unlike existing datasets, SoulDance is a holistic dance dataset that includes body movements, hand gestures, and facial expressions, captured from professional dancers across 284 music segments spanning 15 genres. The data is represented in multiple formats, including skeleton-level BVH files, SMPL-X [41] body models, and FLAME [34] face parameters, providing unprecedented quality and diversity for dance research.
@@ 3 SoulDance Dataset
SoulDance is the first high-quality captured 3D dance dataset that incorporates expressive facial movements in music-dance pairs. As shown in Table 1, most existing datasets lack detailed hand movements and almost entirely omit facial expressions, making it challenging to generate truly holistic dance from such data. To overcome these limitations, we employed a professional holistic motion capture system to record dance movements, ensuring precise synchronization of the body, hands, and face in both temporal and spatial domains, thereby faithfully preserving real-world motion dynamics.
3.1 Data Acquisition
Most dance datasets rely on pose estimation methods [51, 12] to extract dance movements from videos [31, 30, 29]. However, these pipelines often yield low-quality data, as accurate gesture movements and expressions are hard to obtain from video sources alone.
@@ 3.1 Data Acquisition
High-Quality Dance-Music Pairs. We enlisted five renowned professional dancers to perform, ensuring that the dance movements are executed with skill and expertise. Throughout the entire recording process, a director supervised the quality of dance and their alignment with the music, while a engineer was on hand to adjust and correct settings in case of recording misalignment. This rigorous process ensures the capture of high-quality and accurate holistic dance motions. Additionally, we provide an analysis and visual demonstration of the SoulDance dataset. For more details, please refer to Appendix A.
@@ 5.1 Comparison to SOTA Methods
User Study. In the user study, 28 participants (including 7 professional dancers) evaluated 22 video pairs. Results (Table 7) indicate that our SoulDance dataset and SoulNet method outperform existing datasets and methods. See Appendix I for details.
@@ 5.2 Ablation Study
Music-Aligned Strategy. To evaluate the effectiveness of the proposed Music-Aligned strategy, we compare the generation results with and without the MMR module. As shown in Table 5, incorporating the MMR module enhances the quality of the generated dance and improves alignment with the music. Moreover, this enhances the qualitative effects and additional details can be found in Appendix F. This highlights the positive impact of the MMR module on both the quality and synchronization of music-driven dance generation. To further analyze the contributions of each alignment loss, we performed ablation studies on $\mathcal{L}_{\text{Align-body}}$ and $\mathcal{L}_{\text{Align-whole}}$. As shown in Table 6, $\mathcal{L}_{\text{Align-body}}$ significantly improves local feature alignment (FID and BAS), while $\mathcal{L}_{\text{Align-whole}}$ enhances global structure alignment (MMR-MS).
@@ Acknowledgements
We sincerely thank the anonymous ICCV reviewers for their invaluable feedback and suggestions. We would like to thank ByteDance for sponsoring the research. Xiaojie Li would also like to thank his colleagues in the lab, Weixiang Zhang, Rongwei Lu, Jiajun Luo and Duo Wu, for their constructive comments on improving the quality of this paper. This work was supported in part by National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204, 62472249), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001).
@@ Appendix I User Study Details
The duration of the A and B videos ranges from 5 to 8 seconds. The users rated on a scale from 1 to 10, and the scores for each question were denoted as WS, BS, HS, ES, and AS, corresponding to their respective questions in the order mentioned above.

### v2

@@ Abstract
Well-coordinated, music-aligned holistic dance enhances emotional expressiveness and audience engagement. However, generating such dances remains challenging due to the scarcity of holistic 3D dance datasets, the difficulty of achieving cross-modal alignment between music and dance, and the complexity of modeling interdependent motion across the body, hands, and face. To address these challenges, we introduce SoulDance, a high-precision music-dance paired dataset captured via professional motion capture systems, featuring meticulously annotated holistic dance movements. Building on this dataset, we propose SoulNet, a framework designed to generate music-aligned, kinematically coordinated holistic dance sequences. SoulNet consists of three principal components: (1) Hierarchical Residual Vector Quantization, which models complex, fine-grained motion dependencies across the body, hands, and face; (2) Music-Aligned Generative Model, which composes these hierarchical motion units into expressive and coordinated holistic dance; (3) Music-Motion Retrieval Module, a pre-trained cross-modal model that functions as a music-dance alignment prior, ensuring temporal synchronization and semantic coherence between generated dance and input music throughout the generation process. Extensive experiments demonstrate that SoulNet significantly surpasses existing approaches in generating high-quality, music-coordinated, and well-aligned holistic 3D dance sequences. Additional resources are available at: https://xjli360.github.io/SoulDance.
@@ 1 Introduction
Nonetheless, despite recent advances, the generation of well-coordinated and holistic dance sequences that are synchronized with music remains a challenging task, primarily due to three fundamental limitations. The first constraint refers to dataset inadequacies in which existing music-dance datasets either lack crucial components of dance performance or suffer from quality issues. While datasets like AIST++[31] provide body-only movements with musical correspondence, and FineDance[32] and Choreomaster [7] incorporate fine-grained finger movements, they omit facial expressions—a vital element for conveying emotion in dance. Furthermore, most datasets rely on pose estimation from videos, introducing inherent inaccuracies that compromise data quality. The second limitation concerns modeling constraints where current approaches fail to generate coordinated dance movements due to insufficient capacity for capturing complex hierarchical interdependencies between body, hands and face. The third limitation lies in the lack of cross-modal alignment between music and dance, which results in poor synchronization between the two modalities.
To overcome these challenges, we present a comprehensive solution encompassing both dataset and methodological innovations. We introduce SoulDance, a high-quality dataset comprising 12.5 hours of paired music and dance data captured using professional motion capture systems equipped with 15 optical cameras, body motion capture suits, data gloves, and a facial expression tracking system. Unlike existing datasets, SoulDance is a holistic dance dataset that includes body movements, hand gestures, and facial expressions, captured from professional dancers across 284 music segments spanning 15 genres. The data is represented in multiple formats, including skeleton-level BVH files, SMPL-X [41] body models, and FLAME [34] face parameters, providing unprecedented quality and diversity for dance research.
@@ 3 SoulDance Dataset
SoulDance is the first high-quality captured 3D dance dataset that incorporates expressive facial movements in music-dance pairs. As shown in Table 1, most existing datasets lack detailed hand movements and almost entirely omit facial expressions, making it challenging to generate truly holistic dance from such data. Most dance datasets rely on pose estimation methods [51, 12] to extract dance movements from videos [31, 30, 29]. However, these pipelines often yield low-quality data, as accurate gesture movements and expressions are hard to obtain from video sources alone. To overcome these limitations, we employed a professional holistic motion capture system to record dance movements, ensuring precise synchronization of the body, hands, and face in both temporal and spatial domains, thereby faithfully preserving real-world motion dynamics.
@@
High-Quality Dance-Music Pairs. We enlisted five renowned professional dancers to perform, ensuring that the dance movements are executed with skill and expertise. Throughout the entire recording process, a director supervised the quality of dance and their alignment with the music, while a engineer was on hand to adjust and correct settings in case of recording misalignment. Furthermore, expert choreography was incorporated to guarantee perfect alignment between music and dance. This meticulous process resulted in a high-quality, temporally and spatially coherent dance motion-music dataset, providing a solid foundation for expressive and coordinated dance generation. Additionally, we provide an analysis and visual demonstration of the SoulDance dataset. For more details, please refer to Appendix A.
@@ 5.1 Comparison to SOTA Methods
User Study. In the user study, 28 participants—including 7 professional dancers—evaluated 22 pairs of videos. The durations of videos A and B ranged from 5 to 8 seconds. Participants were asked to answer a series of questions (see Appendix I for details), each rated on a scale from 1 to 10. The scores were denoted as Whole Score (WS), Body Score (BS), Hands Score (HS), Emotion Score (ES), and Alignment Score (AS), corresponding to the respective questions in order. The results (Table 7) demonstrate that our proposed SoulDance dataset and SoulNet model outperform existing datasets and methods across all evaluation metrics.
@@ 5.2 Ablation Study
Music-Aligned Strategy. To evaluate the effectiveness of the proposed Music-Aligned strategy, we compare the generation results with and without the MMR module. As shown in Table 5, incorporating the MMR module enhances the quality of the generated dance and improves alignment with the music. Moreover, this enhances the qualitative effects and additional details can be found in Appendix F. To further analyze the contributions of each alignment loss, we performed ablation studies on $\mathcal{L}_{\text{Align-body}}$ and $\mathcal{L}_{\text{Align-whole}}$. As shown in Table 6, $\mathcal{L}_{\text{Align-body}}$ significantly improves local feature alignment (FID and BAS), while $\mathcal{L}_{\text{Align-whole}}$ enhances global structure alignment (MMR-MS).
@@ Acknowledgements
We sincerely thank the anonymous ICCV reviewers for their invaluable feedback and suggestions. We would like to thank ByteDance for sponsoring the research. Xiaojie Li would also like to thank his colleagues in the lab, Weixiang Zhang, Rongwei Lu, Jiajun Luo and Duo Wu, for their constructive comments on improving the quality of this paper. This work was supported in part by National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204, 62472249), National Key R&D Program of China (No. 2022YFF0902303), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001).